OpenStudy (anonymous):

A rope is swinging in such a way that the length of the arc is decreasing geometrically. If the first arc is 18 feet long and the third arc is 8 feet long, what is the length of the second arc? 12 feet 10 feet 5 feet 72 feet

OpenStudy (nurali):

let second rope be x..then x/18 = 8/x or x^2 = 18 x 8 = 144 x =12...this is the answer

OpenStudy (anonymous):

thanks sooooo much! can you help me with another?

OpenStudy (nurali):

My Pleasure i will try.

OpenStudy (anonymous):

okay one sec :) @Nurali

OpenStudy (anonymous):

Use summation notation to write the series 2 + 4 + 6 + 8 + ... for 10 terms.

OpenStudy (anonymous):

thats choice a and at the end its ;2n

OpenStudy (anonymous):

C. n

OpenStudy (nurali):

2+4+.. means its 2 so the answer is 2n

OpenStudy (nurali):

have u got it. @jessicasapimp

OpenStudy (anonymous):

yeah sorta. @Nurali

OpenStudy (anonymous):

so would it be a or d ?

OpenStudy (nurali):

tell me choice a,b,c,d